# The Effect of a Non-universal Extinction Curve on the Wesenheit Function and Cepheid Distances

### v1

@@ 5 Conclusions
Milky Way classical Cepheids are especially affected because they lie in the Galactic disk, where extinction is both high and highly variable. As a result, even moderate changes in $R_{V}$ produce noticeable shifts in the $R$ coefficients used in optical Wesenheit relations. For Gaia-based Wesenheit magnitudes, $W_{G}$, a shift from $R_{V}=3.1$ to $R_{V}=3.6$ leads to changes of $\approx 0.7$ mag in $W_{G}$, corresponding to distance differences of almost 4040%. These shifts are large enough to explain the strong discrepancies found by Skowron2025 between Milky Way Cepheid distances derived from Gaia photometry and those obtained from mid-infrared PL relations combined with extinction maps. Near-infrared Wesenheit indices remain the least affected and therefore offer the most reliable distances.
@@
PW relations constructed with different Wesenheit indices are also widely used for distance measurements in other galaxies, most notably in the Magellanic Clouds. As shown by Zhang2025, the mean $R_{V}$ in the SMC is lower than in the Milky Way, while in the LMC, the range of $R_{V}$ is much wider. Because the choice of $R$ directly affects the definition of the Wesenheit function, adopting an $R_{V}$ appropriate for one galaxy but applying it to another will change the slope and zero point of the PW relation, and therefore alter the inferred distances. This would also affect the calibration of the extragalactic distance scale and, in consequence, influence measurements of the Hubble constant. For example, Riess2016 showed that $R$ variations arising from changing between the reddening laws of Fitzpatrick1999 and Cardelli1989 result in $H_{0}$ changing by 0.10.150.1-0.15 km s-1 Mpc-1 when $W_{H}$ is used, and by 2.153.8-2.15-3.8 km s-1 Mpc-1 when $W_{VI}$ is used, demonstrating that near-infrared data remains the preferred choice for $H_{0}$ studies, in the absence of individual $R$ values.
@@
The main conclusion of this work is that optical Wesenheit indices of Milky Way Cepheids are not ”reddening-free”, as the assumption of a universal extinction curve is not satisfied. However, a positive consequence is that the colors of highly reddened Cepheids contain information of $R_{V}$ on large scales, as will be demonstrated in Drimmel2025. It is also important to note, that although our discussion focuses on classical Cepheids, the same considerations are valid for all classes of pulsating stars, such as RR Lyrae or long-period variables, for which PW relations are employed to derive distances.
6 Software and third party data repository citations
@@
DMS acknowledges support from the European Union (ERC, LSP-MIST, 101040160). SK acknowledges support from the European Union’s Horizon 2020 research and innovation program under the GaiaUnlimited project (grant agreement No 101004110). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. RD & EP are supported in part by the Italian Space Agency (ASI) through contract 2025-10-HH.0 to the National Institute for Astrophysics (INAF). This work made use of the Overleaf platform and the NASA Astrophysics Data System.

### v2

@@ V Conclusions
Milky Way classical Cepheids are especially affected because they lie in the Galactic disk, where extinction is both high and highly variable. As a result, even moderate changes in $R_{V}$ produce noticeable shifts in the $R$ coefficients used in optical Wesenheit relations. For Gaia-based Wesenheit magnitudes, $W_{G}$, a shift from $R_{V}=3.1$ to $R_{V}=3.6$ leads to changes of $\approx 0.7$ mag in $W_{G}$, corresponding to distance differences of almost 4040%. These shifts are large enough to explain the strong discrepancies found by Skowron et al. (2025) between Milky Way Cepheid distances derived from Gaia photometry and those obtained from mid-infrared PL relations combined with extinction maps. Near-infrared Wesenheit indices remain the least affected and therefore offer the most reliable distances.
@@
PW relations constructed with different Wesenheit indices are also widely used for distance measurements in other galaxies, most notably in the Magellanic Clouds. As shown by Zhang and Green (2025), the mean $R_{V}$ in the SMC is lower than in the Milky Way, while in the LMC, the range of $R_{V}$ is much wider. Because the choice of $R$ directly affects the definition of the Wesenheit function, adopting an $R_{V}$ appropriate for one galaxy but applying it to another will change the slope and zero point of the PW relation, and therefore alter the inferred distances. This would also affect the calibration of the extragalactic distance scale and, in consequence, influence measurements of the Hubble constant. For example, Riess et al. (2016) showed that $R$ variations arising from changing between the reddening laws of Fitzpatrick (1999) and Cardelli et al. (1989) result in $H_{0}$ changing by 0.10.150.1-0.15 km s-1 Mpc-1 when $W_{H}$ is used, and by 2.153.8-2.15-3.8 km s-1 Mpc-1 when $W_{VI}$ is used, demonstrating that near-infrared data remains the preferred choice for $H_{0}$ studies, in the absence of individual $R$ values.
@@
The main conclusion of this work is that optical Wesenheit indices of Milky Way Cepheids are not ”reddening-free”, as the assumption of a universal extinction curve is not satisfied. Consequently, optical Wesenheit magnitudes should not be used to estimate distances to Milky Way Cepheids, and distance determinations should instead rely on near- or mid-infrared data, which are far less sensitive to variations in the extinction curve (e.g. Skowron et al. 2025; Wang and Chen 2025). However, a positive consequence is that the colors of highly reddened Cepheids contain information of $R_{V}$ on large scales, as will be demonstrated in Drimmel et al. (2025, in prep.). It is also important to note, that although our discussion focuses on classical Cepheids, the same considerations are valid for all classes of pulsating stars, such as RR Lyrae or long-period variables, for which PW relations are employed to derive distances.
Finally, one could also ask about the continued use of Wesenheit functions in an era when obtaining high-quality multi-band photometry is no longer as big of a problem, as it was four decades ago, when the Wesenheit index was introduced. Multi-band approaches that simultaneously fit an extinction curve to apparent distance moduli across several wavelengths, introduced by Freedman (1988) for $BVRI$ Cepheid photometry in IC 1613, make use of all available information and have been shown to provide reliable estimates of both distance and reddening for many galaxies. This method was later extended by Rich et al. (2014) to a star-by-star basis. They calculated individual distance moduli to each Cepheid in NGC 6822 in seven bands ($VIJHK_{S}$, $3.6$ and $4.5\mu m$), and fitted a reddening law (in distance-modulus versus inverse-wavelength space) for each star to determine the true distance modulus (corresponding to zero inverse wavelength). The Wesenheit index is nothing more than such a fit, but restricted to only two, or three, passbands. It is therefore straightforward that, when multi-band data are available, the Wesenheit function becomes unnecessarily restrictive.
Another important aspect of deriving Cepheid distances is taking into account wavelength-dependent effects such as metallicity, which has been shown to affect the PL and PW relations (e.g. Bhuyan et al. 2026 and references therein). At the same time, metallicity is expected to correlate with dust properties, thus affecting the extinction curve parameters. Therefore, fully accounting for both the effects of variable $R_{V}$ and metallicity, would require a complex covariant approach. While this is beyond the scope of this Letter, it represents an important issue that should be addressed in future studies.
VI Software and third party data repository citations
@@
The authors thank the Referee for insightful and constructive comments, which inspired a broader perspective on the subject of this Letter. DMS acknowledges support from the European Union (ERC, LSP-MIST, 101040160). SK acknowledges support from the European Union’s Horizon 2020 research and innovation program under the GaiaUnlimited project (grant agreement No 101004110). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. RD & EP are supported in part by the Italian Space Agency (ASI) through contract 2025-10-HH.0 to the National Institute for Astrophysics (INAF). This work made use of the Overleaf platform and the NASA Astrophysics Data System.